# Value-oriented forecast reconciliation for renewables in electricity markets

### v1

@@ 4.2 Case 1
We investigate the range of weights in the allocation rule, which vary between 0 and 1, and present the average profit for each WPP across various models in Figure 2. Overall, the profit for each WPP increases as $w$ rises. We have confirmed that the unit balancing cost condition (23) holds in this simulation. As expected by Proposition 3.2, the function $c_{t+k,i}^{(\texttt{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})$ decreases monotonically with increasing $w$, which explains the observed monotonic increase in profits with respect to $w$. Specifically, within the frameworks of both BU reconciliation and quality-oriented reconciliation, profits are directly proportional to $w$. Notably, the quality-oriented reconciliation training is independent of $w$, and similarly, BU reconciliation does not require training. Thus, the balancing cost terms in (22) remain unaffected by changes in $w$, leading to a linear relationship.
@@
Figure 2: Average profit for each wind power producer across different models in Case 1
@@
Additionally, nearly every model results in higher profits for all WPPs, except in the case of quality-focused forecast reconciliation when $w$ equals 00 or $0.1$. Given that the reconciled forecasts at the bottom level remain unchanged in the BU reconciliation framework, it can be inferred that the profits exceed those from independent offers when condition (23) is satisfied. Comparing the profits from BU reconciliation to those from quality-oriented reconciliation reveals that quality-based reconciliation generally yields higher profits, except for WPP $4$. This outcome is attributed to the ability of quality-oriented reconciliation to reduce the overall mean squared errors more effectively than BU reconciliation, as supported by existing research (Van Erven and Cugliari,, 2015). However, the profits for WPP $4$ suggest that quality-oriented reconciliation might lead to an unfair allocation of costs. Specifically, the value-oriented forecast reconciliation generates greater profits compared to other models because its loss function in the reconciliation process aligns with balancing costs.
@@ 4.3 Case 2
Figure 3: Average profit for each wind power producer across different models in Case 2
We then analyze the average profit of each WPP when base forecasts are communicated as quantiles. The analysis considers varying the weight $w$ in the allocation rule from 0 to 1, with the results shown in Figure 3. Similar to Case 1, the average profits for each WPP increase as $w$ rises, although there are a few exceptions under the value-oriented forecast reconciliation approach. Provided that the unit balancing cost condition (23) is satisfied, the profits from BU reconciliation consistently exceed those from independent offering. For the proposed approach, WPP $4$ is an exception when $w=0.1$, primarily due to the challenges in model training at lower values of $w$. In fact, the base forecast profits in this case are substantial, suggesting that the agreement set during bargaining is minimal, which in turn complicates the training process. With more training samples, it is expected to attain estimated parameters within the agreement set. Nevertheless, the profits from the value-oriented forecast reconciliation approach tend to exceed those from BU reconciliation, owing to the information gain discussed in Section 3.2.2.
@@
We compare the average profits resulting from the proposed value-oriented forecast reconciliation approach in two scenarios: when the base forecasts are either mean forecasts or quantile forecasts. The results are presented in Figure 4, where the ‘m’ and ‘q’ in the legend indicate that the base forecasts are mean or quantile. The analysis shows that, when the base forecasts are mean forecasts, the profits generated by the proposed method exceed those of the alternative approach, even though profits under independent offering remain lower. Notably, when $w$ is either $0.9$ or $1.0$, the profits obtained with the proposed method are almost identical in both scenarios. This finding reinforces the earlier discussion by suggesting that the agreement set becomes more restrictive when base forecasts are quantile forecasts, thereby complicating model training for smaller values of $w$. In contrast, this implies that wind power producers need only to provide their mean forecasts to the PM, which reduces their need to analyze the electricity markets.
@@ 5 Real-world Case Study
Quality-oriented reconciliation 15.83 28.27 33.71 23.45
Value-oriented reconciliation 15.92 28.45 33.95 23.62
@@
Similar to Section 4, we also have an hourly forward market followed by a balancing market with a dual-price settlement mechanism. However, unlike the simulation studies, we obtain two years of hourly price data from the Danish TSO data portal [^3] https://www.energidataservice.dk/. The hourly penalties are calculated using the day-ahead spot and regulation prices, as specified in Section (2). We consider the four models described in Section 4. As noted in the previous section, the mean forecasts suffice as the base forecasts; therefore, we define the base forecasts as the mean forecasts. Specifically, the quality-oriented reconciliation model incorporates lagged wind power generation values as contextual features. While for the value-oriented reconciliation model, we include lagged penalties $\psi^{+}_{t-h},\cdots,\psi^{+}_{t}$, $\psi^{-}_{t-h},\cdots,\psi^{-}_{t}$ besides wind power generation values as contextual features. In particular, the weight $w$ in the allocation rule is set to $0.9$.
@@
We present the average profits for each wind power producer in Table 1. The portfolio-based models generate higher average profits compared to independent offering, aligning with the findings from simulation studies. Specifically, BU reconciliation and quality-oriented reconciliation exhibit similar performance, although the latter provides more accurate forecasts. As a result, the profit increase is primarily attributed to participation in aggregate offering. This supports our argument that quality-oriented reconciliation may not be the most effective approach for evaluating forecast accuracy. As expected, value-oriented reconciliation yields the highest average profits for each wind power producer, as it facilitates the development of a profitable trading strategy.
6 Conclusions
@@
Although value is considered a critical factor in evaluating forecasts, alongside quality (Murphy,, 1993), it has been largely overlooked in the forecast reconciliation literature. Generally, forecast reconciliation can be viewed as an information pool, where a group of agents pools their information to make collective decisions. We emphasize the importance of integrating forecast value into the reconciliation process, demonstrating that quality-oriented approaches may fail to produce decisions that are fair and acceptable to all agents. This issue becomes particularly pronounced when agents with heterogeneous loss functions are involved, as quality-oriented reconciliation may lead to individual revenue losses, which in turn can cause potential conflicts. Consequently, we propose a value-oriented forecast reconciliation approach based on the Nash bargaining framework. Specifically, we frame the problem as a learning-based reconciliation problem, where parameters are estimated using a primal-dual algorithm.
@@
We demonstrate the effectiveness of the proposed approach through an aggregated wind energy trading problem, where profits are distributed among wind power producers according to a weighted allocation rule. We introduce the concept of consensus on information, distinguishing it from the well-established concept of consensus on allocations. Specifically, we design simulations and case studies that consider an hourly forward market and a balancing market with a dual-price mechanism, utilizing market data from Denmark and wind power data from the Wind Toolkit. The results show that, under specific conditions, the proposed method ensures that the profits achieved through the portfolio-based approach consistently outperform those obtained through independent offering. This gain results from both information pooling and cooperative trading. Notably, the findings suggest that, in terms of subsequent decision-making value, quality-focused forecast reconciliation may not be the optimal choice.
However, we note that the allocation rule used in this study does not lie within the core of the cooperative trading game. Future investigations should aim to identify efficient allocation rules within the core. Although this paper focuses on point forecast reconciliation, the proposed method can be readily adjusted for information pooling in group decision-making processes. Similar algorithms could also be designed for reconciling probabilistic forecasts and combining forecasts. Another promising avenue for future research is to apply the fairness concept to value-oriented forecasting in a broader context.
@@
$c_{t+k}(\tilde{y}_{t+k,\texttt{sum}|t},y_{t+k,\texttt{sum}})$ $=\psi^{+}_{t+k}\left[y_{t+k,\texttt{sum}}-\tilde{y}_{t+k,\texttt{sum}|t}\right]^{+}+\psi^{-}_{t+k}\left[\tilde{y}_{t+k,\texttt{sum}|t}-y_{t+k,\texttt{sum}}\right]^{+}$
@@
where the inequality results from the sub-additivity property of $[x]^{+}$ for any $x\in\mathbb{R}$. ∎
@@
Using the definition of $c_{t+k,i}^{(\texttt{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})$ and $\gamma_{i}^{\texttt{pc}}$, we have
$c_{t+k,i}^{(\texttt{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})$ $=(1-w)c_{t+k}(\tilde{y}_{t+k,i|t},y_{t+k,i})+w\frac{c_{t+k}(\tilde{y}_{t+k,i|t},y_{t+k,i})c_{t+k}(\tilde{y}_{t+k,\texttt{sum}|t},y_{t+k,\texttt{sum}})}{\sum_{i=1}^{m}c_{t+k}(\tilde{y}_{t+k,i|t},y_{t+k,i})}$
@@
where the inequality results from Theorem 3.1. ∎
@@ Acknowledgment
Honglin Wen is funded by National Science Foundation of China (52307119).

### v2

@@ 4.2 Case 1
Figure 2: Average profit change over independent offers for each wind power producer across different models in Case 1
We investigate the range of weights in the allocation rule, which varies between 0 and 1, and present the average profit over independent offers for each WPP across various models in Figure 2. Overall, each WPP’s profit under independent offerings is invariant to $w$, whereas profits under reconciliation-based strategies increase as $w$ rises. We have confirmed that the unit balancing cost condition (26) holds in this simulation. As expected by Proposition 3.3, the function $c_{t+k,i}^{(\mathrm{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})$ decreases monotonically with increasing $w$, which explains the observed monotonic increase in profits with respect to $w$. Specifically, within the frameworks of both BU reconciliation and quality-oriented reconciliation, profits are directly proportional to $w$. Notably, the quality-oriented reconciliation training is independent of $w$, and similarly, BU reconciliation does not require training. Thus, the balancing cost terms in (25) remain unaffected by changes in $w$, leading to a linear relationship.
Figure 3: Average profit change over independent offers for each wind power producer under learning-based and linear reconciliation in Case 1
@@
Additionally, nearly every reconciliation-based model results in higher profits for all WPPs, except in the case of quality-focused forecast reconciliation when $w$ equals 0 or $0.1$. Given that the reconciled forecasts at the bottom level remain unchanged in the BU reconciliation framework, it can be inferred that the profits exceed those from independent offers when condition (26) is satisfied. Comparing the profits from BU reconciliation to those from quality-oriented reconciliation reveals that quality-based reconciliation generally yields higher profits, except for WPP $4$. This outcome is attributed to the ability of quality-oriented reconciliation to reduce the overall mean squared errors more effectively than BU reconciliation, as supported by existing research (van2015game). However, the profits for WPP $4$ suggest that quality-oriented reconciliation might lead to an unfair allocation of costs. Specifically, the value-oriented forecast reconciliation generates greater profits compared to other models because its loss function in the reconciliation process aligns with balancing costs. Evidently, the gains from aggregated trading differ across WPPs. In particular, WPPs 1 and 2 realize larger percentage profit increases relative to their independent offerings. This is consistent with the Nash product objective, which equalizes and safeguards excess profits over the independent baseline but does not enforce identical percentage gains. Because independent-offer profits differ across WPPs, even identical absolute excess profits translate into different percentage changes.
We present the results under both learning-based and linear reconciliation methods in Figure 3. Overall, value-oriented reconciliation yields higher profits for each WPP compared to quality-oriented reconciliation, supporting the claim that value-oriented reconciliation is agnostic to the choice of reconciliation model. Notably, the performance of quality-oriented reconciliation is similar across both learning-based and linear implementations, highlighting the effectiveness of the optimal reconciliation approach proposed by hyndman2011optimal. In contrast, for value-oriented reconciliation, the learning-based model outperforms its linear counterpart in both performance and stability. This is expected, as value-oriented reconciliation involves a more complex training procedure, and neural networks offer greater flexibility in function approximation.
Figure 4: Average profit change over independent offers for each wind power producer across different models in Case 2
@@ 4.3 Case 2
We then analyze the average profit of each WPP when base forecasts are communicated as quantiles. The analysis considers varying the weight $w$ in the allocation rule from 0 to 1, with the results shown in Figure 4. Similar to Case 1, the average profits for each WPP increase as $w$ rises, although there are a few exceptions under the value-oriented forecast reconciliation approach. Provided that the unit balancing cost condition (26) is satisfied, the profits from BU reconciliation consistently exceed those from independent offering. For quality-oriented reconciliation, the results indicate that the resulting profits are lower than those achieved through strategic independent offering in most cases. This highlights a key limitation of quality-oriented reconciliation when applied in a decision-making context. For the proposed approach, WPP $4$ is an exception when $w=0.1$, primarily due to the challenges in model training at lower values of $w$. In fact, the base forecast profits in this case are substantial, suggesting that the agreement set during bargaining is minimal, which in turn complicates the training process. With more training samples, it is expected to attain estimated parameters within the agreement set. Nevertheless, the profits from the value-oriented forecast reconciliation approach tend to exceed those from BU reconciliation, owing to the information gain discussed in Section 3.2.2.
Table 1: RMSE of hierarchical forecasts in Cases 1 and 2 (MW)
Weight
Case Method 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1.0
1 Reconciliation (value) 0.698 0.685 0.672 0.685 0.681 0.676 0.668 0.668 0.666 0.878
Reconciliation (quality) 0.605 0.598 0.605 0.605 0.599 0.608 0.610 0.610 0.598 0.607
2 Reconciliation (value) 0.689 0.695 0.692 0.689 0.689 0.688 0.682 0.679 0.674 0.806
Reconciliation (quality) 0.601 0.605 0.602 0.604 0.603 0.600 0.608 0.612 0.606 0.602
@@
We compare the absolute average profits resulting from the proposed value-oriented forecast reconciliation approach in both scenarios: when the base forecasts are either mean forecasts or quantile forecasts. The results are presented in Figure 5, where the ‘m’ and ‘q’ in the legend indicate that the base forecasts are mean or quantile. The analysis shows that, when the base forecasts are mean forecasts, the profits generated by the proposed method exceed those of the alternative approach, even though profits under independent offering remain lower. Notably, when $w$ is either $0.9$ or $1.0$, the profits obtained with the proposed method are almost identical in both scenarios. This finding reinforces the earlier discussion by suggesting that the agreement set becomes more restrictive when base forecasts are quantile forecasts, thereby complicating model training for smaller values of $w$. In contrast, this implies that wind power producers need only to provide their mean forecasts to the PM, which reduces their need to analyze the electricity markets. Furthermore, we report the RMSE of hierarchical forecasts for both cases in Tables 1. As observed, the forecast errors under the value-oriented reconciliation approach are consistently higher than those obtained from the quality-oriented reconciliation. When considered alongside the profit analysis, these results support our central claim: minimizing overall forecast errors does not necessarily lead to cost reductions for individual agents. Specifically, when the weight $w$ is set to 1.0, the RMSE under the value-oriented reconciliation approach is relatively high, as the loss function exclusively prioritizes the aggregate level.
@@ 5 Real-world Case Study
Quality-oriented reconciliation $15.83\pm 0.006$ $28.27\pm 0.012$ $33.71\pm 0.016$ $23.45\pm 0.011$
Value-oriented reconciliation $15.92\pm 0.001$ $28.45\pm 0.004$ $33.95\pm 0.003$ $23.62\pm 0.002$
@@
Similar to Section 4, we also have an hourly forward market followed by a balancing market with a dual-price settlement mechanism. However, unlike the simulation studies, we obtain two years of hourly price data from the Danish TSO data portal [^5] https://www.energidataservice.dk/. The hourly penalties are calculated using the day-ahead spot and regulation prices, as specified in Section (2). We consider the four models described in Section 4. As noted in the previous section, the mean forecasts suffice as the base forecasts; therefore, we define the base forecasts as the mean forecasts. Specifically, the quality-oriented reconciliation model incorporates lagged wind power generation values as contextual features. While for the value-oriented reconciliation model, we include lagged penalties $\psi^{+}_{t-h},\cdots,\psi^{+}_{t}$, $\psi^{-}_{t-h},\cdots,\psi^{-}_{t}$ besides wind power generation values as contextual features. In particular, the weight $w$ in the allocation rule is set to $0.9$. For both quality-oriented and value-oriented reconciliation, we repeat the experiments 10 times and report both the mean and standard deviation of the results.
@@
We present the average profits for each wind power producer in Table 2. The aggregate offering models generate higher average profits compared to independent offering, aligning with the findings from simulation studies. Specifically, BU reconciliation and quality-oriented reconciliation exhibit similar performance, although the latter provides more accurate forecasts. As a result, the profit increase is primarily attributed to participation in aggregate offering. This supports our argument that quality-oriented reconciliation may not be the most effective approach for evaluating forecast accuracy. As expected, value-oriented reconciliation yields the highest average profits for each wind power producer, as it facilitates the development of a profitable trading strategy.
6 Concluding Remarks
@@
Although value is considered a critical factor in evaluating forecasts, alongside quality (murphy1993good), it has been largely overlooked in the forecast reconciliation literature. Generally, forecast reconciliation can be viewed as an information pool, where a group of agents pools their information to make collective decisions. We emphasize the importance of integrating forecast value into the reconciliation process, demonstrating that quality-oriented approaches may fail to produce decisions that are fair and acceptable to all agents. This issue becomes particularly pronounced when agents with heterogeneous loss functions are involved, as quality-oriented reconciliation may lead to individual revenue losses, which in turn can cause potential conflicts. Consequently, we propose a value-oriented forecast reconciliation approach based on the Nash bargaining framework. Specifically, we frame the problem as a learning-based reconciliation problem, where parameters are estimated using a primal-dual algorithm.
@@
We introduce the concept of consensus on information, distinguishing it from the well-established notion of consensus on allocations, which serves as a prerequisite for multi-agent cooperation. We demonstrate the effectiveness of the proposed value-oriented reconciliation approach through an aggregated wind energy trading problem, where profits are distributed among wind power producers according to a weighted proportional allocation rule. Furthermore, we prove that by employing this allocation rule, the agreement set of parameters within the Nash bargaining game is guaranteed to be non-empty. Given the estimated reconciliation and allocation mechanism, it is challenging for individual WPPs to explicitly derive the functional form associated with their optimal offers. Instead, our framework aggregates the available information and reconciles the forecasts to determine an optimal offer for each WPP in a data-driven manner. Since the reconciled forecasts are directly used as decisions in the market, this approach may also be interpreted as a form of decision reconciliation. Specifically, we design simulations and case studies that consider an hourly forward market and a balancing market with a dual-price mechanism, utilizing market data from Denmark and wind power data from the Wind Toolkit. The results show that, under specific conditions, the proposed method ensures that the profits achieved through the aggregate offering approach consistently outperform those obtained through independent offering. This gain results from both information pooling and cooperative trading. Notably, the findings suggest that, in terms of subsequent decision-making value, quality-focused forecast reconciliation may not be the optimal choice.
However, we note that the allocation rule employed in this study does not lie within the core of the cooperative trading game. Furthermore, the proposed rule does not ensure that existing agents in the coalition remain unaffected by the inclusion of new members—an essential property referred to as the population monotonic allocation scheme (PMAS) (sprumont1990population).Future work should aim to develop allocation mechanisms that both lie within the core and support coalition expansion without disadvantaging any participant. Indeed, the proposed framework is not limited to energy trading; it is also applicable to various domains in supply chain management where cooperation among agents is essential, such as in retail operations. Although this study focuses primarily on point forecast reconciliation, the proposed framework is readily adaptable to broader group decision-making scenarios involving information pooling. Related algorithms may also be designed for reconciling probabilistic forecasts or aggregating multiple forecast sources. Another promising avenue for future research is the incorporation of fairness principles into value-oriented forecasting in more general and diverse application domains.
@@
$c_{t+k}(\tilde{y}_{t+k,\mathrm{sum}|t},y_{t+k,\mathrm{sum}})$ $=\psi^{+}_{t+k}\left[y_{t+k,\mathrm{sum}}-\tilde{y}_{t+k,\mathrm{sum}|t}\right]^{+}+\psi^{-}_{t+k}\left[\tilde{y}_{t+k,\mathrm{sum}|t}-y_{t+k,\mathrm{sum}}\right]^{+}$
@@
where the inequality results from the sub-additivity property of $x]^{+}$ for any $x\in\mathbb{R}$. ∎
@@
Using the definition of $c_{t+k,i}^{(\mathrm{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})$ and $\gamma_{i}^{\mathrm{pc}}$, we have
$c_{t+k,i}^{(\mathrm{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})$ $=(1-w)c_{t+k}(\tilde{y}_{t+k,i|t},y_{t+k,i})+w\frac{c_{t+k}(\tilde{y}_{t+k,i|t},y_{t+k,i})c_{t+k}(\tilde{y}_{t+k,\mathrm{sum}|t},y_{t+k,\mathrm{sum}})}{\sum_{i=1}^{m}c_{t+k}(\tilde{y}_{t+k,i|t},y_{t+k,i})}$
@@
where the inequality results from Proposition 3.1. ∎
Batch form of the Lagrangian function
Using a batch of samples $\{\hat{\mathbf{y}}_{t+k|t},\mathbf{z}_{t},\mathbf{y}_{t+k}\}_{t\in\mathcal{T}_{b}}$ where $\mathcal{T}_{b}$ represents the set of indices, we write the Lagrangian of batch samples as
$\begin{split}L^{(b)}(\bm{\theta},\bm{\mu})&=-\sum_{i}^{m}\log\left[\frac{1}{|\mathcal{T}_{b}|}\sum_{t\in\mathcal{T}_{b}}\left[c_{t+k}(\hat{y}_{t+k,i|t},y_{t+k,i})-c_{t+k,i}^{(\mathrm{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})\right]\right]\\ &+\sum_{i}^{m}\mu_{i}\left[\frac{1}{|\mathcal{T}_{b}|}\sum_{t\in\mathcal{T}_{b}}\left[c_{t+k,i}^{(\mathrm{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})-c_{t+k}(\hat{y}_{t+k,i|t},y_{t+k,i})\right]\right]^{+},\end{split}$ (32)
where $|\mathcal{T}_{b}|$ represents the cardinality of $\mathcal{T}_{b}$.
@@ Acknowledgment
We are grateful for the detailed and constructive suggestions from three anonymous reviewers and the editor. Honglin Wen is funded by the National Science Foundation of China (52307119). Pierre Pinson acknowledges the support of UKRI through the Global NSF-UKRI Centre EPICS (Electric Power Innovation for a Carbon-free Society Centre – EPICS-UK; grant no. EP/Y025946/1).